# On the Performance of Tri-Hybrid Beamforming Using Pinching Antennas

### v1

@@ I Introduction
Over the past three decades, multiple-input multiple-output (MIMO) technology has fundamentally transformed wireless communications. Recently, MIMO research has entered a new phase that focuses on reconfigurable antennas [1], which complement conventional digital and analog beamforming as an additional adaptive layer. By adjusting their radiation patterns in real time, these antennas enable flexible electromagnetic (EM) beam control and channel reconfiguration. Reconfigurable antennas are now regarded as a key component of the emerging tri-hybrid beamforming architecture, which is expected to play a pivotal role in shaping sixth-generation (6G) wireless networks [2].
Representative examples of reconfigurable antennas for supporting tri-hybrid beamforming include dynamic metasurface antennas, fluid antennas, and movable antennas [3]. Although these approaches provide notable performance gains, their reconfiguration capability is typically confined to apertures spanning only a few to several tens of wavelengths. Such a scale remains insufficient to overcome two fundamental constraints of wave propagation, particularly in the high-frequency bands envisioned for 6G, such as the 7-24 GHz upper mid-bands: severe large-scale free-space path loss and high susceptibility to blockages [4].
@@ III-A3 Power Scaling Law
which, together with [16, Eq. (2.01.18)], leads to the result in (18). This concludes the proof. ∎
@@ IV Numerical Results
We perform numerical simulations to validate the analytical results. Unless stated otherwise, the system parameters are set as follows [7]: carrier frequency $f_{\rm{c}}=28$ GHz, effective refractive index $n_{\rm{eff}}=1.4$, in-waveguide attenuation factor $\kappa=0.08$ dB/m, minimum inter-PA spacing $\Delta_{\min}=\frac{\lambda}{2}$, power budget $P=10$ dBm, and noise power $\sigma^{2}=-90$ dBm. The user is assumed to be uniformly distributed within a rectangular region centered at the origin, with side lengths $D_{x}=50$ m and $D_{y}=20$ m along the $x$- and $y$-axes, respectively. The waveguides are aligned parallel to the $x$-axis at a fixed height $H=3$ m. Their $y$-coordinates are given by $y_{m}=-\frac{D_{y}}{2}+\frac{m-1}{M-1}D_{y}$ for $m\in{\mathcal{M}}$ so that the waveguides uniformly span the entire horizontal coverage area. The $x$-coordinates of the feed points for all waveguides are set to $x_{0}=-\frac{D_{x}}{2}$.

### v2

@@ I Introduction
Over the past three decades, multiple-input multiple-output (MIMO) technology has fundamentally transformed wireless communications. Recently, MIMO research has entered a new phase that focuses on reconfigurable antennas [1], which complement conventional digital and analog beamforming as an additional adaptive layer. By adjusting their radiation patterns in real time, these antennas enable flexible electromagnetic (EM) beam control and channel reconfiguration. Reconfigurable antennas are now regarded as a key component of the emerging tri-hybrid beamforming architecture, which is expected to play a pivotal role in shaping sixth-generation (6G) wireless networks.
Representative examples of reconfigurable antennas for supporting tri-hybrid beamforming include dynamic metasurface antennas, fluid antennas, and movable antennas [2, 3, 4]. Although these approaches provide notable performance gains, their reconfiguration capability is typically confined to apertures spanning only a few to several tens of wavelengths. Such a scale remains insufficient to overcome two fundamental constraints of wave propagation, particularly in the high-frequency bands envisioned for 6G, such as the 7-24 GHz upper mid-bands: severe large-scale free-space path loss and high susceptibility to blockages [5].
@@ III-A3 Power Scaling Law
which, together with [18, Eq. (2.01.18)], leads to the result in (18). This concludes the proof. ∎
@@ III-B Multiple-RF Scenario
III-C Extension
In the above discussion, the performance of a single user under a TDMA framework was considered. It was assumed that pinching beamforming is redesigned in each time slot for the scheduled user so that the signals radiated by different PAs combine constructively at the intended receiver. This strategy achieves the performance upper bound, but it incurs relatively high implementation cost, since the PA locations must be adjusted in every time slot according to the scheduled user’s location. This operating mode is referred to as pinching switching (PS) [20]. An alternative approach that reduces implementation cost is pinching multiplexing (PM), where all users share a common pinching beamformer across different time slots, i.e., the PA placement remains fixed. Under this scheme, the achievable user rate is generally reduced, but the system implementation becomes significantly simpler because PA repositioning across time slots is avoided.
For simplicity, it is assumed that the digital and analog beamformers of each user are redesigned within its dedicated time slot. These beamformers can be updated electronically with relatively low overhead, whereas adjusting PA positions may require mechanical operations and therefore incurs higher latency and complexity. Under this assumption, the achievable sum-rate of PM-based TDMA can be formulated as follows:
${\mathcal{R}}_{\rm{sum}}\triangleq\frac{1}{K}\sum\nolimits_{k=1}^{K}\log_{2}(1+{P}\lVert\hat{\mathbf{h}}_{k}^{T}{\mathbf{G}}\rVert^{2}/{\sigma^{2}}),$ (32)
where $P$ denotes the per-user transmit power budget, $\sigma^{2}$ denotes the noise power, and $\hat{\mathbf{h}}_{k}\in{\mathbb{C}}^{MN\times 1}$ represents the channel vector from all pinched waveguides to user $k$. Following (10), the effective channel gain $a_{k}({\mathbf{W}}_{\rm{pin}})\triangleq\lVert\hat{\mathbf{h}}_{k}^{T}{\mathbf{G}}\rVert^{2}$ for each user $k$ can be written as follows:
$a_{k}({\mathbf{W}}_{\rm{pin}})=\sum_{m=1}^{M}\left\lvert\sum_{n=1}^{N}\frac{\sqrt{\eta}{\rm{e}}^{-{\rm{j}}\frac{2\pi}{\lambda}r_{m,n}^{(k)}-{\rm{j}}\frac{2\pi}{\lambda_{\rm{g}}}x_{m,n}}}{\sqrt{N}r_{m,n}^{(k)}}\right\rvert$, (33)
where ${\mathbf{W}}_{\rm{pin}}]_{m,n}=x_{m,n}$ for $m\in{\mathcal{M}}$ and $n\in{\mathcal{N}}$. The position of user $k$ is denoted by ${\mathbf{r}}_{k}=[x_{{\rm{u}}}^{(k)},y_{{\rm{u}}}^{(k)},0]^{T}\in{\mathbb{R}}^{3\times 1}$, and $r_{m,n}^{(k)}\triangleq\lVert{\bm{\psi}}_{m,n}-{\mathbf{r}}_{k}\rVert=\sqrt{(x_{m,n}-x_{{\rm{u}}}^{(k)})^{2}+H_{k,m}^{2}}$, with $H_{k,m}\triangleq\sqrt{(y_{m}-y_{{\rm{u}}}^{(k)})^{2}+H^{2}}$ representing the effective elevation distance between user $k$ and the $m$th waveguide. Accordingly, the sum-rate maximization problem under PM-based TDMA can be formulated as follows:
$\max_{{\mathbf{W}}_{\rm{pin}}}$~ ${\mathcal{R}}_{\rm{sum}}=\sum\nolimits_{k=1}^{K}\log_{2}(1+{P}a_{k}({\mathbf{W}}_{\rm{pin}})/{\sigma^{2}})$ (34a)
[…]
Finding the optimal pinching beamformer for PM is more challenging than for PS, since all users share the same pinching beamformer ${\mathbf{W}}_{\rm{pin}}$. As a result, ${\mathbf{W}}_{\rm{pin}}$ must be designed such that the signals intended for different users can be constructively combined at their respective receivers across different time slots, which generally leads to a difficult nonconvex optimization problem. As a compromise between performance and complexity, we employ an element-wise alternating optimization framework to efficiently obtain a high-quality suboptimal solution to the non-convex problem in (34).
In the proposed approach, each antenna position $x_{m,n}$ is optimized sequentially while keeping all other coordinates fixed. Let ${\mathcal{R}}_{m,n}(x)$ denotes the sum-rate when $x_{m,n}$ in ${\mathbf{W}}_{\rm{pin}}$ is set to $x$ while others are fixed. This subproblem is equivalent to the following:
[…]
Since this is a single-variable optimization over a bounded interval, it can be solved efficiently via one-dimensional search. To do so, we discretize the interval $x_{0},x_{\max}]$ into a uniform grid of $X$ points:
${\mathcal{X}}\triangleq\left\{x_{0},x_{0}+\frac{I_{\rm{P}}}{X-1},x_{0}+\frac{2I_{\rm{P}}}{X-1},\ldots,x_{\max}\right\},$ (36)
where $I_{\rm{P}}\triangleq x_{\max}-x_{0}$. A near-optimal $x_{m,n}$, denoted as $x_{m,n}^{\star}$, is then selected according to the following:
$x_{m,n}^{\star}\triangleq\operatorname*{argmax}\nolimits_{x_{m,n}\in{\mathcal{X}}\setminus{\mathcal{X}}_{m,n}}{\mathcal{R}}_{m,n}(x),$ (37)
where the set ${\mathcal{X}}_{m,n}$ includes all grid points that violate the minimum spacing constraint, i.e.,
${\mathcal{X}}_{m,n}\triangleq\{x|x\in{\mathcal{X}},\lvert x-x_{m,n^{\prime}}\rvert<\Delta,n^{\prime}\neq n\}.$ (38)
This procedure is applied iteratively to all antennas until convergence. The complete algorithm is summarized in Algorithm 1, with a computational complexity of ${\mathcal{O}}(I_{\rm{iter}}NMX)$, where $I_{\rm{iter}}$ is the number of iterations to convergence.
Algorithm 1 Element-wise Algorithm for Solving (34)
1: initialize the optimization variables
2:repeat
3:  for $n\in{\mathcal{N}}$ and $m\in{\mathcal{M}}$ do
4:   update $x_{m,n}$ through one-dimensional search
5:  end for
6:until convergence
@@ IV Numerical Results
We perform numerical simulations to validate the analytical results. Unless stated otherwise, the system parameters are set as follows [8]: carrier frequency $f_{\rm{c}}=28$ GHz, effective refractive index $n_{\rm{eff}}=1.4$, in-waveguide attenuation factor $\kappa=0.08$ dB/m, minimum inter-PA spacing $\Delta_{\min}=\frac{\lambda}{2}$, power budget $P=10$ dBm, and noise power $\sigma^{2}=-90$ dBm. The users are assumed to be uniformly distributed within a rectangular region centered at the origin, with side lengths $D_{x}=50$ m and $D_{y}=20$ m along the $x$- and $y$-axes, respectively. The waveguides are aligned parallel to the $x$-axis at a fixed height $H=3$ m. Their $y$-coordinates are given by $y_{m}=-\frac{D_{y}}{2}+\frac{m-1}{M-1}D_{y}$ for $m\in{\mathcal{M}}$ so that the waveguides uniformly span the entire horizontal coverage area. The $x$-coordinates of the feed points for all waveguides are set to $x_{0}=-\frac{D_{x}}{2}$.
@@
Figure 6: Sum-rate versus the user number. $M=4$ and $N=32$.
Figure 6 compares the sum-rate achieved by PS and PM as a function of the number of users. It can be observed that the sum rate achieved by PS remains nearly constant as the number of users increases. This behavior arises because the pinching beamformer is optimized independently in each time slot for the scheduled user, which allows constructive signal combining tailored to each user’s location. In contrast, under PM, all users share the same pinching beamformer. As the number of users increases, this shared beamformer becomes increasingly suboptimal for individual users, which causes the achievable sum rate to decrease. Moreover, the in-waveguide propagation loss under PM becomes more pronounced. This effect occurs because the pinching antennas must be distributed more broadly along the waveguide in order to serve users at different locations, which increases the average propagation distance within the waveguide and exacerbates attenuation.